OpenStudy (anonymous):

How to plug in \[ log_5 (10) \] in calculator

OpenStudy (saifoo.khan):

You may use the change of base rule.

OpenStudy (anonymous):

which is?

OpenStudy (saifoo.khan):

\[\log_5 10 \to \frac{\log 10}{\log 5}\]

OpenStudy (anonymous):

what saifoo said

OpenStudy (anonymous):

Are there other rules?

OpenStudy (anonymous):

the golden rule is a good one

OpenStudy (anonymous):

for log, that is... Are there other rules for log's besides laws

OpenStudy (anonymous):

tons

OpenStudy (anonymous):

Hmm, okay. Thanks everyone!!